# Erwin: A Tree-based Hierarchical Transformer for Large-scale Physical Systems

### v1

@@ Appendix C Experimental details
We follow the experimental setup of the benchmark. The training was done for $5000$ epochs with batch size $16$ for point transformers and batch size $8$ for message-passing-based models. The implementation of SEGNN, NequIP and MPNN was done in JAX and taken from the original benchmark repository (Balla et al., 2024). We maintained the hyperparameters of the baselines used in the benchmark. For Erwin and PointTransformer, those are provided in Table 5. In Erwin’s embedding, we conditioned messages on Bessel basis functions rather than the relative position, which significantly improved overall performance.
@@
All models were trained with batch size $32$ for $50000$ training iterations with an initial learning rate of $5\cdot 10^{-4}$. We finetuned the hyperparameters of every model on the validation dataset (reported in Table 7).
@@
Baseline results are taken from (Janny et al., 2023), except for runtime and peak memory usage, which we measured ourselves. Erwin was trained with batch size $12$ for $4000$ epochs.
Airflow pressure modeling
We take the results of baseline models from Alkin et al. (2024a). Both Erwin and PointTransformer v3 were trained with batch size $32$ for $1000$ epochs, and their hyperparameters are given in Table 6).
@@
Point Grid size 0.01
Transformer v3 Enc. depths (2, 2, 6, 2)
@@
Point Grid size 0.01
Transformer v3 Enc. depths (2, 2, 2, 2, 2)
@@
Point Grid size 0.025
Transformer v3 Enc. depths (2, 2, 2, 6, 2)

### v2

@@ B.7 Further details per experiment
We follow the experimental setup of the benchmark. The training was done for 5,000 epochs with batch size 16 for point transformers and batch size 8 for message-passing-based models. The implementation of SEGNN, NequIP, and MPNN was done in JAX and taken from the original benchmark repository (Balla et al., 2024). We maintained the hyperparameters of the baselines used in the benchmark. For Erwin and PTv3, the hyperparameters are provided in Table 9. In Erwin’s embedding, we conditioned messages on Bessel basis functions rather than the relative position, which significantly improved overall performance.
@@
All models were trained with batch size 32 for 50,000 training iterations with an initial learning rate of $5\cdot 10^{-4}$. We fine-tuned the hyperparameters of every model on the validation dataset (reported in Table 11).
PDE benchmarks
The baseline results and experimental setups are taken from Luo et al. (2025). We adjusted batch size, ball size, and the number of attention heads per block for the best performance. The hidden dimensionality of Erwin was adjusted such that the overall number of parameters is around $10^{6}$, which is comparable with other baselines. Constrained by the parameter size, the same configuration worked the best; see Table 12 for details.
Airflow pressure modeling
We take the results of baseline models from Bleeker et al. (2025). Both Erwin and PTv3 were trained with batch size 32 for 1,000 epochs, and their hyperparameters are given in Table 10. We tuned the number of blocks, the number of message-passing steps, hidden dimensionality, and ball size per block for the best performance. When experimenting with pooling, we found that not involving any coarsening significantly improves model performance; hence, we used stride 1 in each block.
@@
Baseline results are taken from (Janny et al., 2023), except for runtime and peak memory usage, which we measured ourselves. Erwin was trained with batch size 12 for 4,000 epochs. We tuned hidden dimensionality and ball size per block for the lowest validation loss.
@@
PTv3 Grid size 0.01
Enc. depths (2, 2, 6, 2)
@@
PTv3 Grid size 0.01
Enc. depths (2, 2, 2, 2, 2)
@@
PTv3 Grid size 0.025
Enc. depths (2, 2, 2, 6, 2)
@@
Table 12: Model architectures for the PDE benchmarks.
Model Parameter Value
Erwin MPNN dim. 64
Channels 64
Window size 256
Enc. heads (8, 8)
Enc. depths (6, 6)
Dec. heads (8, 8)
Dec. depths (6)
Pooling (1, 1)